# Hierarchical certification of non-classical network correlations

### v3

@@ 2.1 The tripartite chain
2.2 Long chains
The bilocality inequality can be extended to chain networks of arbitrary length by using the notion of $k$-independent sets of parties [31], and with it building appropriate operators $I$ and $J$ in order to construct Bell-like inequalities whose violations witness $\ell$-NN and $\ell$-QNN. Indeed, Ref. [31] showed that a suitable generalisation of the $I$ and $J$ operators is
$I=\Pi_{{\rm{}odd}\,i}A^{+}_{x_{i}}\Pi_{{\rm{}even}\,j}A_{x_{j}=0},\qquad\qquad J=\Pi_{{\rm{}odd}\,i}A^{-}_{x_{i}}\Pi_{{\rm{}even}\,j}A_{x_{j}=1},$ (10)
with $A^{\pm}_{x_{i}}=(A_{x_{i}=0}\pm A_{x_{i}=1})/2$, and the correlators representing the combination of probabilities given by $A_{x_{1}}\cdots A_{x_{n}}=\sum_{\vec{a}}(-1)^{\sum_{i=1}^{n}a_{i}}P(\vec{a}|\vec{x})$.
In the following we analyze the case where the network contains an odd number of parties and the non-classical sources are require to follow quantum theory (i.e., we restrict to the case of $\ell$-QNN in chain networks with an odd number of parties). We defer the analysis of the remaining cases, namely calculations for $\ell$-NN and for chain networks with an even number of parties, to Appendix A. We furthermore distinguish two sub-cases, namely when the network contains $\ell\geq(n-1)/2$ classical sources, and when it contains $\ell<(n-1)/2$ classical sources.
Case $\mathbf{\ell\geq(n-1)/2}$: In this case, there are more classical than quantum sources. This implies that there are no more than $(n-1)/2$ non-adjacent pairs of parties $(\textsf{A}_{i},\textsf{A}_{i+1})$ who share a quantum system, and thus, the network is a $k$-independent quantum network [31] with $k\leq(n-1)/2$. Let $S$ consist of all odd integers, $i$, such that $\textsf{A}_{i}$ shares classical variables with both $\textsf{A}_{i-1}$ and $\textsf{A}_{i+1}$. We have $|S|\geq 1$, and from the classicality of the measurements of $\textsf{A}_{i}$ when $i\in S$ it follows that
$I^{\frac{2}{n+1}}+J^{\frac{2}{n+1}}$ $=\left[\prod\limits_{{\rm{}odd}\,i}A^{+}_{x_{i}}\prod\limits_{{\rm{}even}\,j}A_{x_{j}=0}\right]^{\frac{2}{n+1}}+\left[\prod\limits_{{\rm{}odd}\,i}A^{-}_{x_{i}}\prod\limits_{{\rm{}even}\,j}A_{x_{j}=1}\right]^{\frac{2}{n+1}}$
$\leq\prod\limits_{i\in S}(A^{+}_{x_{i}})^{\frac{2}{n+1}}\left[\prod\limits_{{\rm{}odd}\,i;i\not\in S}A^{+}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=0}\right]^{\frac{2}{n+1}}$
$\quad+\prod\limits_{i\in S}(A^{-}_{x_{i}})^{\frac{2}{n+1}}\left[\prod\limits_{{\rm{}odd}\,i;i\not\in S}A^{-}_{x_{i}}\prod\limits_{{\rm{}even}\,j}A_{x_{j}=1}\right]^{\frac{2}{n+1}},$ (11)
@@ 2.2 Long chains
From the quantum bound for the $\frac{n-1}{2}$-independent quantum network (i.e., the case when no pair of $\textsf{A}_{i}$ with even $i$ share a quantum source) the following inequality holds [31]:
$\left[\prod\limits_{{\rm odd}\,i;i\not\in S}A^{+}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=0}\right]^{\frac{2}{n-1}}+\left[\prod\limits_{{\rm odd}\,i;i\not\in S}A^{-}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=1}\right]^{\frac{2}{n-1}}\leq\sqrt{2}.$ (12)
Considering this, and taking into account that $0\leq[\Pi_{{\rm odd}\,i;i\not\in S}A^{+}_{x_{i}}\Pi_{{\rm even}\,j}A_{x_{j}=0}]^{\frac{2}{n-1}}\leq 1$ and that $0\leq[\Pi_{{\rm odd}\,i;i\not\in S}A^{-}_{x_{i}}\Pi_{{\rm even}\,j}A_{x_{j}=1}]^{\frac{2}{n-1}}\leq 1$, it is possible to write that
$\left[\prod\limits_{{\rm odd}\,i;i\not\in S}A^{+}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=0}\right]^{\frac{2}{n-1}}\leq\sqrt{2}\cos^{2}\theta,$ (13)
$\left[\prod\limits_{{\rm odd}\,i;i\not\in S}A^{-}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=1}\right]^{\frac{2}{n-1}}\leq\sqrt{2}\sin^{2}\theta,$
for some $\theta\in[0,\pi/2]$. Moreover, using the fact that $|A_{x_{i}}|\leq 1$ when $i\in S$ it follows that
$\left|\prod\limits_{i\in S}A^{+}_{x_{i}}\pm\prod\limits_{i\in S}A^{-}_{x_{i}}\right|\leq\,$ $\frac{1}{2^{|S|-1}}\max\left\{\left|\sum_{(x_{i},i\in S);\oplus_{i\in S}x_{i}=0}\prod_{i\in S}A_{x_{i}}\right|,\left|\sum_{(x_{i},i\in S);\oplus_{i\in S}x_{i}=1}\prod_{i\in S}A_{x_{i}}\right|\right\}$
[…]
=\,= $1,$
because |$|\{(x_{i},i\in S)|\oplus_{i\in S}x_{i}=0\}|=|\{(x_{i},i\in S)|\oplus_{i\in S}x_{i}=1\}|=2^{|S|-1}$, the operation $\oplus$ representing addition modulo 2. It is thus possible to write $|\prod\limits_{i\in S}A^{+}_{x_{i}}|\leq\cos^{2}\vartheta$ and $|\prod\limits_{i\in S}A^{-}_{x_{i}}|\leq\sin^{2}\vartheta$ for some $\vartheta\in[0,\pi/2]$. Substituting all the above in Eq. (11) it follows that
$I^{\frac{2}{n+1}}+J^{\frac{2}{n+1}}\leq\,$ $2^{\frac{n-1}{2(n+1)}}\left[\prod\limits_{i\in S}(A^{+}_{x_{i}})^{\frac{2}{n+1}}\cos^{\frac{2n-2}{n+1}}\theta+\prod\limits_{i\in S}(A^{-}_{x_{i}})^{\frac{2}{n+1}}\sin^{\frac{2n-2}{n+1}}\theta\right]$
[…]
@@
Case $\mathbf{\ell<(n-1)/2}$: An example of this is, e.g., when the sources connecting parties $\textsf{A}_{2i}$ and $\textsf{A}_{2i+1}$ distribute classical variables, and the sources connecting parties $\textsf{A}_{2i-1}$ and $\textsf{A}_{2i}$ distribute quantum systems (recall that $i=1,\dots,(n+1)/2$. This network, and in fact any network with fewer than half of its sources being classical, is a $(n+1)/2$-independent quantum network [31], and so the joint correlations satisfy the inequality
@@
$I^{\frac{2}{n+1}}+J^{\frac{2}{n+1}}\leq\begin{cases}2^{\frac{n-1}{2(n+1)}}&\ell\geq\frac{n-1}{2}\\ \sqrt{2}&\ell<\frac{n-1}{2}\end{cases}.$ (17)
There are a few aspects to note of this inequality. First, the bounds obtained are valid for any $\ell$ within each of the ranges, and thus, a violation guarantees the strongest case (namely, any value above $2^{\frac{n-1}{2(n+1)}}$ guarantees $\frac{n-1}{2}$-QNN and any value above $\sqrt{2}$ guarantees FQNN). Second, the bound for $\ell\geq\frac{n-1}{2}$ increases with $n$, which means that it is more demanding to demonstrate $\ell$-QNN with increasing $n$. Third, the gap between the two bounds closes in the limit of $n\rightarrow\infty$. Thus, in infinite chain networks demonstrating non-classicality of even just one of the sources is as demanding as demonstrating non-classicality of all of them. Finally, $\sqrt{2}$ is the maximum value achievable with quantum sources [31], so Eq. (17) cannot be used to witness FQNN. In Appendix A we derive analogous inequalities in the case when the chain has an even number of parties, and that are satisfied by $\ell$-NL correlations.
3 Nonlinear $\ell$-(Q)NN hierarchies for star networks
@@
Figure 3: The $n+1$-partite star network. Correlations that do not admit a realization neither in the form of (a) nor with any other positioning of the classical variable are FQNN, and those that do not admit a realization neither in the form of (b) nor with any other positioning of the classical variable are FNN. There exist correlations of both types that can be produced in the setup depicted in (c).
Let us begin with the analysis of $\ell$-QNN for fixed $\ell$. In this case, we set that the central party can perform two measurements $B_{y}$, $y\in\{0,1\}$, with binary outcomes. $b\in\{0,1\}$. In this setup, the quantities $I$ and $J$ are given by $I=\Pi_{i=1}^{n}A^{+}_{x_{i}}B_{0}$ and $J=\Pi_{i=1}^{n}A^{-}_{x_{i}}B_{1}$ [26]. Consider thus the network in Fig. 3(a), where there are $\ell$ parties sharing classical variables with B while the remaining ones share quantum states. Without loss of generality, one can assume that said parties are $\textsf{A}_{1},\dots,\textsf{A}_{\ell}$. It follows that
$|I|^{\frac{1}{n}}+|J|^{\frac{1}{n}}=\left|\prod\limits_{i=1}^{\ell}A^{+}_{x_{i}}\right|^{\frac{1}{n}}\left|\left(\prod\limits_{j=\ell+1}^{n}A^{+}_{x_{j}}\right)B_{0}\right|^{\frac{1}{n}}+\,\left|\prod\limits_{i=1}^{\ell}A^{-}_{x_{i}}\right|^{\frac{1}{n}}\left|\left(\prod\limits_{j=\ell+1}^{n}A^{-}_{x_{j}}\right)B_{1}\right|^{\frac{1}{n}},$ (18)
@@
$\left|\left(\prod\limits_{j=\ell+1}^{n}A^{+}_{x_{j}}\right)B_{0}\right|^{\frac{1}{n-\ell}}+\left|\left(\prod\limits_{j=\ell+1}^{n}A^{-}_{x_{j}}\right)B_{1}\right|^{\frac{1}{n-\ell}}\leq\sqrt{2}.$
@@
$\left|\left(\prod\limits_{j=\ell+1}^{n}A^{+}_{x_{j}}\right)B_{0}\right|^{\frac{1}{n-\ell}}\leq\sqrt{2}\cos^{2}\theta,\qquad\left|\left(\prod\limits_{j=\ell+1}^{n}A^{-}_{x_{j}}\right)B_{1}\right|^{\frac{1}{n-\ell}}\leq\sqrt{2}\sin^{2}\theta,$ (19)
for some $\theta\in[0,\pi/2]$. Additionally, since all the parties $\textsf{A}_{1},\dots,\textsf{A}_{\ell}$ are independent, we have the classical bounds $|\prod_{i=1}^{\ell}A^{\pm}_{x_{i}}|\leq\prod_{i=1}^{\ell}|A^{\pm}_{x_{i}}|$. In combination with $|A^{+}_{x_{i}}\pm A^{-}_{x_{i}}|=|A_{x_{i}}|\leq 1$, one can define $\prod_{i=1}^{\ell}A^{+}_{x_{i}}\leq\prod_{i=1}^{\ell}\cos^{2}\vartheta_{i}$ and $\prod_{i=1}^{\ell}A^{-}_{x_{i}}\leq\prod_{i=1}^{\ell}\sin^{2}\vartheta_{i}$ for some $\vartheta_{i}\in[0,\pi/2]$. Substituting this in Eq. (18) we have that
$|I|^{\frac{1}{n}}+|J|^{\frac{1}{n}}=2^{\frac{n-\ell}{2n}}\left(\prod_{i=1}^{\ell}\cos^{\frac{2}{n}}\vartheta_{i}\cos^{\frac{2n-2\ell}{n}}\theta+\prod_{i=1}^{\ell}\sin^{\frac{2}{n}}\vartheta_{i}\sin^{\frac{2n-2\ell}{n}}\theta\right).$
The maximum of the right-hand side is achieved at $\vartheta_{1}=\cdots=\vartheta_{\ell}:=\vartheta$. This implies that
$|I|^{\frac{1}{n}}+|J|^{\frac{1}{n}}\leq 2^{\frac{n-\ell}{2n}}\left(\cos^{\frac{2\ell}{n}}\vartheta\cos^{\frac{2n-2\ell}{n}}\theta+\sin^{\frac{2\ell}{n}}\vartheta\sin^{\frac{2n-2\ell}{n}}\theta\right).$
@@
The right-hand side of (20) has a different value for each $\ell$, thereby giving a hierarchy of inequalities that detects $\ell$-QNN for every $\ell\in[1,\cdots,m]$.
Consider now the situation when there are $\ell$ extremal parties (e.g., $\textsf{A}_{1},\dots,\textsf{A}_{\ell}$) that share classical randomness with B while the remaining parties share non-signaling systems. In this case, the bounds in Eq. (19) are replaced by $|\left(\Pi_{i>\ell}A^{+}_{x_{i}}\right)B_{0}|\leq 1$, $|\left(\Pi_{i>\ell}A^{-}_{x_{i}}\right)B_{1}|\leq 1$, since there exist distributions built out of PR boxes that saturate these bounds, in analogy with the construction in Appendix A.2 for chain networks. We thus have that
$|I|^{\frac{1}{n}}+|J|^{\frac{1}{n}}\leq\,$ $\left|\prod\limits_{i=1}^{\ell}A^{+}_{x_{i}}\right|^{\frac{1}{n}}+\left|\prod\limits_{i=1}^{\ell}A^{-}_{x_{i}}\right|^{\frac{1}{n}}$
@@
where we have used the fact that $|A^{+}_{x_{i}}|\leq\cos^{2}\vartheta_{i}$ and $|A^{-}_{x_{i}}|\leq\sin^{2}\vartheta_{i}$ for some $\vartheta_{i}$ in Eq. (21), which stem from the inequality $|A^{+}_{x_{i}}\pm A^{-}_{x_{i}}|\leq 1$. Inequality (22) follows in a way analogous to that of $\ell$-QNN studied above, and Eq. (23) follows from computing the unique extreme point of the continuous function $\cos^{\frac{2\ell}{n}}\vartheta+\sin^{\frac{2\ell}{n}}\vartheta$, which is $\vartheta=\pi/4$. We thus, again, obtain a hierarchy of bounds that detects $\ell$-NN for every $\ell$. Note also that, for any placement of the classical sources, the same inequality is obtained, and thus a violation of the single inequality (23) is a certification of $\ell$-NN.
Both sets of inequalities, Eqs. (20), (23), can be violated in quantum mechanics. If each of the sources distributes a bipartite quantum system in the state given by Eq. (9), and the parties perform the measurements given by $A_{x_{i}}=\cos\vartheta\sigma_{3}+(-1)^{x_{i}}\sin\vartheta\sigma_{1}$ and $B_{y}=(1-y)\sigma_{3}^{\otimes n}+y\sigma_{1}^{\otimes n}$ [31] one has that

### v4

@@ III.1 The tripartite chain
III.2 Long chains
The bilocality inequality can be extended to chain networks of arbitrary length by using the notion of $k$-independent sets of parties [31], [^1] As defined in Ref. [31], $k$-independence indicates that the maximum number of parties in the network that share no source with each other is $k$. and with it building appropriate operators $I$ and $J$ in order to construct Bell-like inequalities whose violations witness $\ell$-NN and $\ell$-QNN. Indeed, Ref. [31] showed that a suitable generalization of the $I$ and $J$ operators is
$I$ $=\left\langle\Pi_{{\rm{}odd}\,i}A^{+}_{x_{i}}\Pi_{{\rm{}even}\,j}A_{x_{j}=0}\right\rangle,$
$J$ $=\left\langle\Pi_{{\rm{}odd}\,i}A^{-}_{x_{i}}\Pi_{{\rm{}even}\,j}A_{x_{j}=1}\right\rangle,$
with $A^{\pm}_{x_{i}}=(A_{x_{i}=0}\pm A_{x_{i}=1})/2$, and the correlators representing the combination of probabilities given by $\left\langle A_{x_{1}}\cdots A_{x_{n}}\right\rangle=\sum_{\vec{a}}(-1)^{\sum_{i=1}^{n}a_{i}}P(\vec{a}|\vec{x})$. Reference [31] proved that for any $k$-independent network the correlations generated by classical sources satisfy the Bell inequality $|I|^{\frac{1}{k}}+|J|^{\frac{1}{k}}\leq 1$, and that this inequality admits quantum violations using maximally entangled and GHZ states.
In the following we analyze the case where the network contains an odd number of parties and the nonclassical sources are required to follow quantum theory. In other words, we restrict to the case of $\ell$-QNN in chain networks with an odd number of parties. We defer the analysis of the remaining cases, namely calculations for $\ell$-NN and for chain networks with an even number of parties, to Appendix A. We furthermore distinguish two sub-cases, namely when the network contains $\ell\geq(n-1)/2$ classical sources, and when it contains $\ell<(n-1)/2$ classical sources.
Case $\mathbf{\ell\geq(n-1)/2}$: In this case, there are more classical than quantum sources. This implies that there are no more than $(n-1)/2$ non-adjacent pairs of parties $(\textsf{A}_{i},\textsf{A}_{i+1})$ who share a quantum system, and thus, the network is a $k$-independent network [31] with $k\leq(n-1)/2$. Let $S$ consist of all odd integers, $i$, such that $\textsf{A}_{i}$ receives only classical sources. The fact that there are more classical than quantum sources implies that $|S|\geq 1$, and from the classicality of the measurements of $\textsf{A}_{i}$ when $i\in S$ it follows that
$I^{\frac{2}{n+1}}$ $+J^{\frac{2}{n+1}}=\left\langle\prod\limits_{{\rm{}odd}\,i}A^{+}_{x_{i}}\prod\limits_{{\rm{}even}\,j}A_{x_{j}=0}\right\rangle^{\frac{2}{n+1}}$
$\qquad\qquad\quad+\left\langle\prod\limits_{{\rm{}odd}\,i}A^{-}_{x_{i}}\prod\limits_{{\rm{}even}\,j}A_{x_{j}=1}\right\rangle^{\frac{2}{n+1}}$
$\leq\left\langle\prod\limits_{i\in S}A^{+}_{x_{i}}\right\rangle^{\frac{2}{n+1}}\left\langle\prod\limits_{{\rm{}odd}\,i;i\not\in S}A^{+}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=0}\right\rangle^{\frac{2}{n+1}}$
$\quad+\left\langle\prod\limits_{i\in S}A^{-}_{x_{i}}\right\rangle^{\frac{2}{n+1}}\left\langle\prod\limits_{{\rm{}odd}\,i;i\not\in S}A^{-}_{x_{i}}\prod\limits_{{\rm{}even}\,j}A_{x_{j}=1}\right\rangle^{\frac{2}{n+1}}\!\!\!\!\!,$ (9)
@@ III.2 Long chains
Reference [31] gave a quantum bound in the $k$-independent network, which when $k=\frac{n-1}{2}$ reads:
$\left\langle\prod\limits_{{\rm odd}\,i;i\not\in S}A^{+}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=0}\right\rangle^{\frac{2}{n-1}}$
$+\left\langle\prod\limits_{{\rm odd}\,i;i\not\in S}A^{-}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=1}\right\rangle^{\frac{2}{n-1}}\leq\sqrt{2}.$
Considering this, and taking into account that $0\leq\left\langle\Pi_{{\rm odd}\,i;i\not\in S}A^{+}_{x_{i}}\Pi_{{\rm even}\,j}A_{x_{j}=0}\right\rangle^{\frac{2}{n-1}}\leq 1$ and that $0\leq\left\langle\Pi_{{\rm odd}\,i;i\not\in S}A^{-}_{x_{i}}\Pi_{{\rm even}\,j}A_{x_{j}=1}\right\rangle^{\frac{2}{n-1}}\leq 1$, it is possible to write that
$\left\langle\prod\limits_{{\rm odd}\,i;i\not\in S}A^{+}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=0}\right\rangle^{\frac{2}{n-1}}\leq\sqrt{2}\cos^{2}\theta,$ (10)
$\left\langle\prod\limits_{{\rm odd}\,i;i\not\in S}A^{-}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=1}\right\rangle^{\frac{2}{n-1}}\leq\sqrt{2}\sin^{2}\theta,$
for some $\theta\in[0,\pi/2]$. Moreover, using the fact that $|\left\langle A_{x_{i}}\right\rangle|\leq 1$ when $i\in S$ it follows that
$\left|\left\langle\prod\limits_{i\in S}A^{+}_{x_{i}}\pm\prod\limits_{i\in S}A^{-}_{x_{i}}\right\rangle\right|$
$\qquad\leq\frac{1}{2^{|S|-1}}\max\left\{\left|\sum_{(x_{i},i\in S);\oplus_{i\in S}x_{i}=0}\left\langle\prod_{i\in S}A_{x_{i}}\right\rangle\right|\right.,$
$\qquad\qquad\qquad\qquad\quad\left.\left|\sum_{(x_{i},i\in S);\oplus_{i\in S}x_{i}=1}\left\langle\prod_{i\in S}A_{x_{i}}\right\rangle\right|\right\}$
$\qquad\leq\frac{1}{2^{|S|-1}}\times 2^{|S|-1}$
$\qquad=1,$
because |$|\{(x_{i},i\in S)|\oplus_{i\in S}x_{i}=0\}|=|\{(x_{i},i\in S)|\oplus_{i\in S}x_{i}=1\}|=2^{|S|-1}$, the operation $\oplus$ representing addition modulo 2. It is thus possible to write $\left|\left\langle\prod\limits_{i\in S}A^{+}_{x_{i}}\right\rangle\right|\leq\cos^{2}\vartheta$ and $\left|\left\langle\prod\limits_{i\in S}A^{-}_{x_{i}}\right\rangle\right|\leq\sin^{2}\vartheta$ for some $\vartheta\in[0,\pi/2]$. Substituting all the above in Eq. (9) it follows that
$I^{\frac{2}{n+1}}+J^{\frac{2}{n+1}}\leq$ $2^{\frac{n-1}{2(n+1)}}\left[\left\langle\prod\limits_{i\in S}A^{+}_{x_{i}}\right\rangle^{\frac{2}{n+1}}\cos^{\frac{2n-2}{n+1}}\theta\right.$ (11)
$\qquad\qquad+\left.\left\langle\prod\limits_{i\in S}A^{-}_{x_{i}}\right\rangle^{\frac{2}{n+1}}\sin^{\frac{2n-2}{n+1}}\theta\right]$
$\leq$ $2^{\frac{n-1}{2(n+1)}}\left(\cos^{\frac{4}{n+1}}\vartheta\cos^{\frac{2n-2}{n+1}}\theta\right.$
$\qquad\qquad+\left.\sin^{\frac{4}{n+1}}\vartheta\sin^{\frac{2n-2}{n+1}}\theta\right).$
@@
Case $\mathbf{\ell<(n-1)/2}$: An example of this is, e.g., when the sources connecting parties $\textsf{A}_{2i}$ and $\textsf{A}_{2i+1}$ distribute classical variables, and the sources connecting parties $\textsf{A}_{2i-1}$ and $\textsf{A}_{2i}$ distribute quantum systems (recall that $i=1,\dots,(n+1)/2$). This network, and in fact any network with fewer than half of its sources being classical, is a $(n+1)/2$-independent network [31], and so the joint correlations satisfy the inequality
@@
$I^{\frac{2}{n+1}}+J^{\frac{2}{n+1}}\leq\begin{cases}2^{\frac{n-1}{2(n+1)}}&\ell\geq\frac{n-1}{2}\\ \sqrt{2}&\ell<\frac{n-1}{2}.\end{cases}$ (14)
There are a few aspects to note of this inequality. First, the bounds obtained are valid for any $\ell$ within each of the ranges, and thus, a violation guarantees the strongest case (namely, any value above $2^{\frac{n-1}{2(n+1)}}$ guarantees $\frac{n-1}{2}$-QNN and any value above $\sqrt{2}$ guarantees FQNN). Second, the bound for $\ell\geq\frac{n-1}{2}$ increases with $n$, which means that it is more demanding to demonstrate $\ell$-QNN with increasing $n$. Third, the gap between the two bounds closes in the limit of $n\rightarrow\infty$. Thus, in infinite chain networks demonstrating nonclassicality of even just one of the sources is as demanding as demonstrating nonclassicality of all of them. Finally, $\sqrt{2}$ is the maximum value achievable with quantum sources [31], so Eq. (14) cannot be used to witness FQNN. In Appendix A we derive analogous inequalities in the case when the chain has an even number of parties, and that are satisfied by $\ell$-NL correlations.
IV Nonlinear $\ell$-(Q)NN hierarchies for star networks
@@
Figure 3: The $n+1$-partite star network. Correlations that do not admit a realization neither in the form of panel (a) nor with any other positioning of the classical variable are FQNN, and those that do not admit a realization neither in the form of panel (b) nor with any other positioning of the classical variable are FNN. There exist correlations of both types that can be produced in the setup depicted in panel (c).
Let us begin with the analysis of $\ell$-QNN for fixed $\ell$. In this case, we set that the central party can perform two measurements $B_{y}$, $y\in\{0,1\}$, with binary outcomes. $b\in\{0,1\}$. In this setup, the quantities $I$ and $J$ are given by $I=\left\langle B_{0}\Pi_{i=1}^{n}A^{+}_{x_{i}}\right\rangle$ and $J=\left\langle B_{1}\Pi_{i=1}^{n}A^{-}_{x_{i}}\right\rangle$ [26]. Consider thus the network in Fig. 3(a), where there are $\ell$ parties sharing classical variables with B while the remaining ones share quantum states. Without loss of generality, one can assume that said parties are $\textsf{A}_{1},\dots,\textsf{A}_{\ell}$. It follows that
$|I|^{\frac{1}{n}}+|J|^{\frac{1}{n}}=$ $\left|\left\langle\prod\limits_{i=1}^{\ell}A^{+}_{x_{i}}\right\rangle\right|^{\frac{1}{n}}\left|\left\langle B_{0}\prod\limits_{j=\ell+1}^{n}A^{+}_{x_{j}}\right\rangle\right|^{\frac{1}{n}}$
$+\,\left|\left\langle\prod\limits_{i=1}^{\ell}A^{-}_{x_{i}}\right\rangle\right|^{\frac{1}{n}}\left|\left\langle B_{1}\prod\limits_{j=\ell+1}^{n}A^{-}_{x_{j}}\right\rangle\right|^{\frac{1}{n}},$ (15)
@@
$\left|\left\langle B_{0}\prod\limits_{j=\ell+1}^{n}A^{+}_{x_{j}}\right\rangle\right|^{\frac{1}{n-\ell}}+\left|\left\langle B_{1}\prod\limits_{j=\ell+1}^{n}A^{-}_{x_{j}}\right\rangle\right|^{\frac{1}{n-\ell}}\leq\sqrt{2}.$
@@
$\left|\left\langle B_{0}\prod\limits_{j=\ell+1}^{n}A^{+}_{x_{j}}\right\rangle\right|^{\frac{1}{n-\ell}}\leq\sqrt{2}\cos^{2}\theta,$ (16)
$\left|\left\langle B_{1}\prod\limits_{j=\ell+1}^{n}A^{-}_{x_{j}}\right\rangle\right|^{\frac{1}{n-\ell}}\leq\sqrt{2}\sin^{2}\theta,$
for some $\theta\in[0,\pi/2]$. Additionally, since all the parties $\textsf{A}_{1},\dots,\textsf{A}_{\ell}$ are independent, we have the classical bounds $\left|\left\langle\prod_{i=1}^{\ell}A^{\pm}_{x_{i}}\right\rangle\right|\leq\prod_{i=1}^{\ell}\left|\left\langle A^{\pm}_{x_{i}}\right\rangle\right|$. In combination with $\left|\left\langle A^{+}_{x_{i}}\pm A^{-}_{x_{i}}\right\rangle\right|=\left|\left\langle A_{x_{i}}\right\rangle\right|\leq 1$, one can define $\left|\left\langle\prod_{i=1}^{\ell}A^{+}_{x_{i}}\right\rangle\right|\leq\prod_{i=1}^{\ell}\cos^{2}\vartheta_{i}$ and $\left|\left\langle\prod_{i=1}^{\ell}A^{-}_{x_{i}}\right\rangle\right|\leq\prod_{i=1}^{\ell}\sin^{2}\vartheta_{i}$ for some $\vartheta_{i}\in[0,\pi/2]$. Substituting this in Eq. (15) we have that
$|I|^{\frac{1}{n}}+|J|^{\frac{1}{n}}=2^{\frac{n-\ell}{2n}}$ $\left(\prod_{i=1}^{\ell}\cos^{\frac{2}{n}}\vartheta_{i}\cos^{\frac{2n-2\ell}{n}}\theta\right.$
$\quad\left.+\prod_{i=1}^{\ell}\sin^{\frac{2}{n}}\vartheta_{i}\sin^{\frac{2n-2\ell}{n}}\theta\right).$
The maximum of the right-hand side is achieved at $\vartheta_{1}=\dots=\vartheta_{\ell}:=\vartheta$. This implies that
$|I|^{\frac{1}{n}}+|J|^{\frac{1}{n}}\!\leq\!2^{\frac{n-\ell}{2n}}\!\!\left(\cos^{\frac{2\ell}{n}}\vartheta\cos^{\frac{2n-2\ell}{n}}\theta\!+\!\sin^{\frac{2\ell}{n}}\vartheta\sin^{\frac{2n-2\ell}{n}}\theta\right)\!.$
@@
The right-hand side of Eq. (17) has a different value for each $\ell$, thereby giving a hierarchy of inequalities that detects $\ell$-QNN for every $\ell\in[1,\cdots,m]$.
Consider now the situation when there are $\ell$ extremal parties (e.g., $\textsf{A}_{1},\dots,\textsf{A}_{\ell}$) that share classical randomness with B while the remaining parties share nonsignaling systems. In this case, the bounds in Eq. (16) are replaced by $\left|\left\langle B_{0}\Pi_{i>\ell}A^{+}_{x_{i}}\right\rangle\right|\leq 1$, $\left|\left\langle B_{1}\Pi_{i>\ell}A^{-}_{x_{i}}\right\rangle\right|\leq 1$, since there exist distributions built out of PR boxes that saturate these bounds, in analogy with the construction in Appendix A.2 for chain networks. We thus have that
$|I|^{\frac{1}{n}}+|J|^{\frac{1}{n}}\leq\,$ $\left|\left\langle\prod\limits_{i=1}^{\ell}A^{+}_{x_{i}}\right\rangle\right|^{\frac{1}{n}}+\left|\left\langle\prod\limits_{i=1}^{\ell}A^{-}_{x_{i}}\right\rangle\right|^{\frac{1}{n}}$
@@
where we have used the fact that $\left|\left\langle A^{+}_{x_{i}}\right\rangle\right|\leq\cos^{2}\vartheta_{i}$ and $\left|\left\langle A^{-}_{x_{i}}\right\rangle\right|\leq\sin^{2}\vartheta_{i}$ for some $\vartheta_{i}$ in Eq. (18), which stem from the inequality $\left|\left\langle A^{+}_{x_{i}}\pm A^{-}_{x_{i}}\right\rangle\right|\leq 1$. Inequality (19) follows in a way analogous to that of $\ell$-QNN studied above, and Eq. (20) follows from computing the unique extreme point of the continuous function $\cos^{\frac{2\ell}{n}}\vartheta+\sin^{\frac{2\ell}{n}}\vartheta$, which is $\vartheta=\pi/4$. We thus, again, obtain a hierarchy of bounds that detects $\ell$-NN for every $\ell$. Note also that, for any placement of the classical sources, the same inequality is obtained, and thus a violation of the single inequality (20) is a certification of $\ell$-NN.
Both sets of inequalities, Eqs. (17) and (20), can be violated in quantum mechanics. If each of the sources distributes a bipartite quantum system in the state given by Eq. (8), and the parties perform the measurements given by $A_{x_{i}}=\cos\vartheta\sigma_{3}+(-1)^{x_{i}}\sin\vartheta\sigma_{1}$ and $B_{y}=(1-y)\sigma_{3}^{\otimes n}+y\sigma_{1}^{\otimes n}$ [31] one has that